OpenStudy (anonymous):

What do these look like?? Associative Property of Multiplication Distributive Property Associative Property of Addition Commutative Property of Addition

OpenStudy (anonymous):

Associative property of multiplication: a(bc) = (ab)c e.g. (2 * 3)*5 = 6*5 = 30 = 2 * (3*5) = 2* 15 Distributive property: a(b+c) = ab + ac e.g. 2(3+5) = 2(8) = 16 = 2x3 + 2x5 = 6 + 10 = 16 Associative property of addition (a+b) + c =a + (b+c) e.g. (2+3) + 5 = 5 + 5 = 10 = 2 + (3+5) = 2 + 8 Commutative property of addition: a + b = b + a

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

No problem